OpenStudy (anonymous):

Determine the distance between P(3 , -2 , 5) and the plane 2x + 4y - z = 2. Please show all work. Thanks (:

OpenStudy (anonymous):

2x + 4y - z = 2 this is plane equation in R3 (2,4,-1) it's the perpendicular vector to this plane the corresponding unit vector is (2,4,-1)/sqrt21 the point P(3 , -2 , 5) is since 2*3+4*(-2)+5*(-1)=6-8-5=-7 not equals 2 distance is -7-2/sqrt21=-9/sqrt21

OpenStudy (anonymous):

So the answer is : -7-2/sqrt21=-9/sqrt21 =-1.96?

OpenStudy (anonymous):

would be positive. The - sign means that is on the side of the plane which doen's contain the origin. But i don't think you need that. So just use +

OpenStudy (anonymous):

But how did you even solve up to that point? I dont understand what you did

OpenStudy (anonymous):

....

OpenStudy (anonymous):

 i explain you for the case of the line in R2. In R3 is the same. n is the unit normal vector to the line and r is the position vector. So the equation of the line in vector form is: r.n=c, where c is constant. What this equation states, that the projection on the unit normal of the position vector of the points belonging to the line is constant. You can folow the rest from here

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

What is R2 and R3?

OpenStudy (anonymous):

spaces: 2 dimention, 3 dimention of the real numbers pairs, triplets

OpenStudy (anonymous):

I have never dealt with this, I am in 12 Calculus

OpenStudy (anonymous):

no idea what is 12 Calculus

OpenStudy (anonymous):

Grade 12 Calculus

OpenStudy (anonymous):

you mean school?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

highschool

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, more easy way:

OpenStudy (anonymous):

find any point Q on the plane. It means 3 numbers(x,y,z) matching the 2x + 4y - z = 2 for example (0,0,-2). Take vector PQ and do dot product with perpendicular vector to the plane, (2,4,-1) devided by magnitude of perpendicular vector , sqrt21. This is the distanece

OpenStudy (anonymous):

How do you get 21?

OpenStudy (anonymous):

if v(x,y,z) is a vector, it's length is: \[|v|=\sqrt{x ^{2}+y ^{2}+z ^{2}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you take me step-by-step on how you get the answer?

OpenStudy (anonymous):

Im trying to pretend its like a test question, so my teacher would mark my work

OpenStudy (anonymous):

**work shown

OpenStudy (anonymous):

PQ=(3,-2,7) (3,-2,7).(2,4,-1)=6-8-7=-9 deviding by sqrt21 you get -9/sqrt21

OpenStudy (anonymous):

So if I show that work, I would get full marks?

OpenStudy (anonymous):

i guess, if this is the way you supose to do it, :)

OpenStudy (anonymous):

no idea about highschool way...

OpenStudy (anonymous):

What methods are you using to solve this question?

OpenStudy (anonymous):

length of vector dot product

OpenStudy (anonymous):

These are the chapter titles: Cartesion vectors, dot product, Applications of the dot product, Vectors in three space, the cross product and its properties and applications of the dot prosuct and cross product

OpenStudy (anonymous):

so then this is the way

OpenStudy (anonymous):

What are the methods you used?

OpenStudy (anonymous):

PQ=(3,-2,7) isnt this supposed to be (3, -2, 5)?

OpenStudy (anonymous):